OpenStudy (anonymous):

In a classroom, 8 out of 19 students surveyed say they had two or more siblings. If the entire school has 570 students, predict the number of students that have two or more siblings. (Points : 1) 330 240 30 11

OpenStudy (mathstudent55):

Use a proportion. 8 is to 19 as x is to 570. Can you write that as a proportion?

OpenStudy (anonymous):

I don't get it and proportion

OpenStudy (mathstudent55):

A proportion is two ratios that are equal. You are given 8 out of 19. Can you write that as a ratio?

OpenStudy (anonymous):

8:19 correct

OpenStudy (mathstudent55):

Good. 8:19 is the same as 8/19 and \(\dfrac{8}{19} \) and "8 to 19." These are all correct ways of writing a ratio.

OpenStudy (anonymous):

oh ok

OpenStudy (mathstudent55):

The way of writing a ratio that is more practical for us is when you write is as a fraction. 8 out of 19 menas \( \dfrac{8}{19} \) We already have one ratio. Now we need to find another ratio and set the two ratios equal. That is what is called a proportion.

OpenStudy (anonymous):

oh got it

OpenStudy (mathstudent55):

Since the school has 570 students, and we want to figure out how many of those students have two or more siblings, we will call that number x, since we don't know it yet.

OpenStudy (mathstudent55):

Just like in the classroom the ratio is 8 out of 19, the ratio for the whole school is x out of 570.

OpenStudy (anonymous):

ooh ok

OpenStudy (mathstudent55):

Now we write x out of 570 as a ratio in fraction form. That is \(\dfrac{x}{570} \), and we have our second ratio. Now we set the two ratios equal: \( \dfrac{8}{19} = \dfrac{x}{570} \) All that's left is to solve for x.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

To solve a proportion, one easy method is called "cross multiplication." That means multiply the numerator of one fraction by the denominator of the other fraction across and set that equal to other numerator and denominator multiplied together. They make a cross (or x) pattern; that's why it's called cross multiplication. See the figure below.

OpenStudy (mathstudent55):